# Transversity Generalized Parton Distributions of $\Delta$ with the Diquark Spectator Model

### v1

@@ 3.1 Numerical transversity GPDs and helicity amplitudes
the different quark GPDs in all the $\Delta$ isobars can be obtained just by counting the corresponding quark number. With the approximation $c_{2}=c_{3}=0$, only $H^{qT}_{1,3,5,7}$, all of which respectively corresponds to the Lorentz structure of spin-1/2, survive and other transversity GPDs vanish as explained in Sec. 2.2. The non-zero transversity GPDs are shown in Fig. 4 as the functions of variables $x$ and $t$ with different skewness $\xi$, in which $\xi=(0,-0.2,-0.4)$, respectively. The constraint $|\xi|\leq 1/\sqrt{1-4M^{2}/t}$ leads to a bound on the squared momentum transfer $t\geq\frac{4M^{2}\xi^{2}}{1-\xi^{2}}$. Due to time reversal constraints in Eq. (5), only the negative $\xi$ region is shown in the results. We have verified the sum rules of the GPDs in Eq. (10), $\int_{-1}^{1}\mathrm{d}xH^{qT}_{3}(x,\xi,t)=0$. Similar with unpolarized and longitudinal polarized GPDs [48], there is also a tendency that the maximums or minimums of $\xi$-even transversity GPDs $H^{qT}_{1,5,7}$ shift to large $x$ as $|\xi|$ increases. In order to illustrate this character more intuitively and to observe the change with different $\xi$, the 2D cutting planes of $H_{1}^{qT}$ with different $\xi$ and the corresponding minimal t-t are listed in Fig. 5.
Figure 5: (a): The GPDs $H^{qT}_{1}(x,\xi,-|t|_{\min})$ with different skewness $\xi$ and the corresponding minimal $|t|$. The red solid line, orange dashed line and blue dot dash line respectively represent the GPDs with $\left(\xi,-t\right)=\left(0,0\,\text{GeV}^{2}\right)$, $\left(-0.2,0.2\,\text{GeV}^{2}\right)$ and $\left(-0.4,1.0\,\text{GeV}^{2}\right)$. (b): The transversity PDF $h_{1}(x)$, where $h_{1}(x)=2H_{1}^{qT}(x,0,0)$.
Because of the integral instability around the boundary, we take a larger $t=1.0\,\text{GeV}^{2}$ instead of $0.9\,\text{GeV}^{2}$ when $\xi=-0.4$. One can assume that the ratio of the parton momentum fraction equals the mass ratio, ii.ee. $\frac{x+|\xi|}{1-x}=\frac{m_{q}}{m_{D}}$, and this assumption determines a position $x_{\text{max}}=\frac{m_{q}+m_{D}|\xi|}{m_{q}+m_{D}}$. The positions $x_{\text{max}}=0.345,\,0.476,\,0.607$ respectively corresponds to $\xi=0,\,-0.2,\,-0.4$. Furthermore, one can find that the $\xi$-even GPDs are concentrated near $x_{\text{max}}$. This character implies that the parton momenta are distributed by their masses and it agrees with our intuition. Moreover, the GPDs corresponding to the unpolarized and longitudinaly-polarized PDFs also have this feature [48].
@@
where the labels - and ++ are the emitted and re-absorbed quark helicities, respectively, and the explicit forms of the helicity-flip amplitudes are given in Ref. [47]. According to the explicit forms in Ref. [47], the helicity amplitudes can be expressed as $\mathcal{A}^{q}_{\lambda^{\prime}-,\lambda+}=F(\zeta)\mathcal{A}^{\prime q}_{\lambda^{\prime}-,\lambda+}$, where $F(\zeta)$ is the complex part and $\mathcal{A}^{\prime}_{\lambda^{\prime}-,\lambda+}$ is the real part. The specific form of the complex part is $F(\zeta)=C^{\zeta}\theta(\zeta)+C^{*-\zeta}\theta(-\zeta)$ with $\zeta=\lambda^{\prime}-\lambda+1$, where $C^{\zeta}$ represents the $\zeta$ powers of $C$ and $\theta(\zeta)$ is the Heaviside step function with $\theta(0)=\frac{1}{2}$ and $F(0)=1$. And $F(\zeta)$ will be real when the helicity conserve, $\lambda^{\prime}-\lambda=-1$. Moreover, the factor $C$ carries all the complex phase information,

### v2

@@ 3.1 Numerical transversity GPDs and helicity amplitudes
the different quark GPDs in all the $\Delta$ isobars can be obtained just by counting the corresponding quark number. With the approximation $c_{2}=c_{3}=0$, only $H^{qT}_{1,3,5,7}$, all of which respectively corresponds to the Lorentz structure of spin-1/2, survive and other transversity GPDs vanish as explained in Sec. 2.2. The non-zero transversity GPDs are shown in Fig. 4 as the functions of variables $x$ and $t$ with different skewness $\xi$, in which $\xi=(0,-0.2,-0.4)$, respectively. The constraint $|\xi|\leq 1/\sqrt{1-4M^{2}/t}$ leads to a bound on the squared momentum transfer $t\geq\frac{4M^{2}\xi^{2}}{1-\xi^{2}}$, from which one can get the corresponding $|t|_{\text{min}}\sim(0,-0.2,-0.9)$. Due to time reversal constraints in Eq. (5), only the negative $\xi$ region is shown in the results. We have verified the sum rules of the GPDs in Eq. (10), $\int_{-1}^{1}\mathrm{d}xH^{qT}_{3}(x,\xi,t)=0$. Similar with unpolarized and longitudinal polarized GPDs [48], there is also a tendency that the maximums or minimums of $\xi$-even transversity GPDs $H^{qT}_{1,5,7}$ shift to large $x$ as $|\xi|$ increases. In order to illustrate this character more intuitively and to observe the change with different $\xi$, the 2D cutting planes of $H_{1}^{qT}$ with different $\xi$ and the corresponding minimal t-t are listed in Fig. 5.
Figure 5: (a): The GPDs $H^{qT}_{1}(x,\xi,-|t|_{\min})$ with different skewness $\xi$ and the corresponding minimal $|t|$. The red solid line, orange dashed line and blue dot dash line respectively represent the GPDs with $\left(\xi,-t\right)=\left(0,0\,\text{GeV}^{2}\right)$, $\left(-0.2,0.2\,\text{GeV}^{2}\right)$ and $\left(-0.4,0.9\,\text{GeV}^{2}\right)$. (b): The transversity PDF $h_{1}(x)$, where $h_{1}(x)=2H_{1}^{qT}(x,0,0)$.
One can assume that the ratio of the parton momentum fraction equals the mass ratio, ii.ee. $\frac{x+|\xi|}{1-x}=\frac{m_{q}}{m_{D}}$, and this assumption determines a position $x_{\text{max}}=\frac{m_{q}+m_{D}|\xi|}{m_{q}+m_{D}}$. The positions $x_{\text{max}}=0.345,\,0.476,\,0.607$ respectively corresponds to $\xi=0,\,-0.2,\,-0.4$. Furthermore, one can find that the $\xi$-even GPDs are concentrated near $x_{\text{max}}$. This character implies that the parton momenta are distributed by their masses and it agrees with our intuition. Moreover, the GPDs corresponding to the unpolarized and longitudinaly-polarized PDFs also have this feature [48].
@@
where the labels - and ++ are the emitted and re-absorbed quark helicities, respectively, and the explicit forms of the helicity-flip amplitudes are given in Ref. [47]. According to the explicit forms in Ref. [47], the helicity amplitudes can be expressed as $\mathcal{A}^{q}_{\lambda^{\prime}-,\lambda+}=F(\zeta)\mathcal{A}^{\prime q}_{\lambda^{\prime}-,\lambda+}$, where $F(\zeta)$ is the complex part and $\mathcal{A}^{\prime}_{\lambda^{\prime}-,\lambda+}$ is the real part. The specific form of the complex part is $F(\zeta)=C^{\zeta}\theta(\zeta)+C^{*-\zeta}\theta(-\zeta)$ with $\zeta=\lambda^{\prime}-\lambda+1$, where $C^{\zeta}$ represents the $\zeta$ powers of $C$ and $\theta(\zeta)$ is the Heaviside step function with $\theta(0)=\frac{1}{2}$ and $F(0)=1$. The factor $F(\zeta)$ will be real when the helicity conserves, $\lambda^{\prime}-\lambda=-1$ (i.e. the helicity conservation implies that the helicity transfer of the hadron corresponds to that of the parton). Moreover, the factor $C$ carries all the complex phase information,